OpenStudy (anonymous):

Change 4.98 radians to degree measure. Round to the nearest tenth.

OpenStudy (primeralph):

Multiply by 180/pi.

OpenStudy (praxer):

1 radian = 180/ pi degree. so, 4.98 radian = 180/pi x 4.98 degree @raijuenea1 can u do the rest :)